# ROGR: Relightable 3D Objects using Generative Relighting

### v1

@@ 1 Introduction
Inserting real-world objects into new environments is a long-standing problem in computer graphics [1, 2], with numerous applications in the movie and gaming industries. While recent years have seen significant progress in 3D object reconstruction for view synthesis using radiance fields [3, 4], these techniques represent an object illuminated by a single fixed environment and they do not enable changing the appearance of the object due to changes in the lighting. This work extends 3D object reconstruction to enable rendering the object under arbitrary illumination.
@@
At the same time, recent relighting diffusion models [5, 6, 7] have demonstrated impressive capabilities for generating realistic images of objects under arbitrary illumination. However, they only generate a single relit image at a time, which results in inconsistent relighting results when applied to a sequence of viewpoints. While these inconsistently-relit samples can be reconstructed into a single 3D representation [7], optimizing a new 3D representation for each new target lighting is tedious and precludes interactive use cases.
@@ 2 Related Work
Recovering relightable 3D representations of objects from images is a longstanding goal in computer vision and graphics. The prevalent approach is inverse rendering: decomposing the object’s appearance into its underlying geometry, illumination, and material parameters, and relighting the object by simulating the physical interaction of a new target illumination with the recovered geometry and materials [8, 9, 10, 11].
Modern methods for reconstructing relightable 3D representations with inverse rendering use differentiable rendering techniques [12, 13] to optimize mesh [14, 15], distance field [16], or volumetric [17, 18, 19, 20, 21, 22, 23, 24] representations of object geometry and material parameters. While these inverse rendering approaches can be effective, they come with significant limitations: errors in estimating an object’s geometry and materials can produce unrealistic appearance when the object is relit under a new illumination and physically-accurate relighting involves computationally-expensive Monte Carlo simulation of global illumination, which can be too slow for interactive use cases.
@@
Early approaches for interactive relighting in the field of computer graphics proposed the idea of Precomputed Radiance Transfer (PRT) [25]. As appearance behaves linearly with respect to lighting, an object or scene’s appearance can be precomputed under a set of basis lightings and these can be linearly combined to produce appearance under any desired target lighting condition. While PRT-based techniques enable interactive relighting, they struggle with the memory demands of storing the precomputed transfer matrices. To enable rendering under novel illuminations from arbitrary viewpoints, PRT-based methods need to store a full transfer matrix for each 3D point in the scene. These memory requirements are prohibitive even for modestly sized scenes and environment maps, so a large body of work focuses on compressing these PRT matrices [26, 27].
One-light-at-a-time (OLAT) captures in light stages [28] can be thought of as directly capturing basis vectors of a real object’s radiance transfer matrices. Since each image captured by the light stage is of the object illuminated by a single element of a standard lighting basis, they can be linearly combined to reproduce the object’s appearance under any desired environment illumination.
@@
In the deep learning era, many methods have trained networks to directly output relit images. Early techniques utilized standard convolutional neural networks [29, 30] and more recent approaches are based on powerful diffusion models [5, 7, 6, 31]. These generative relighting models have produced impressive results for image relighting, but they cannot be easily used for 3D relighting as the relit appearance is often not consistent across views. IllumiNeRF [7], MIS [31], and Neural Gaffer [6] perform 3D relighting by reconciling samples from a 2D image relighting diffusion model into a single NeRF. However, the relit 3D representation needs to be optimized separately for each novel target illumination. Recent work on 3D reconstruction from images of an object with varying illumination [32] uses a multi-view diffusion model to relight them to have consistent illumination. However, this approach does not allow relighting using a new unseen illumination condition, so it does not allow for generalization. And it does not utilize the high-frequency details in environment maps. Zeng et al. [33] and RNG [34] utilize shadow and highlight cues as conditioning signals for NeRF. In contrast, we explicitly use the encoded features of incident light from the specular reflection direction. Additionally, in contrast to all of these approaches, our proposed method trains a generalizable relightable NeRF that can be rendered under any target illumination without additional optimization. Concurrent work RelitLRM [35] recovers geometry and appearance but fails to maintain consistent geometry across illuminations. DiffusionRenderer [36] uses video diffusion for relighting but lacks 3D consistency. In contrast, our method enforces consistent geometry under varying lighting via explicit modeling and rendering, while also enabling fast inference without the long sampling times of standard diffusion.
@@ 3.1 Multi-View Diffusion Relighting
Our work crucially relies on a multi-view diffusion relighting model to provide consistent relit images, which can then be distilled into a relightable 3D model. This is in contrast with prior work on relighting using diffusion models [7, 6] which independently relights single images and results in ambiguities that must be resolved at the 3D reconstruction stage.
Our network architecture is inspired by multi-view diffusion architectures such as CAT3D [37], which start with pretrained 2D image diffusion models and inflate them by adding cross-attention layers to process multiple views. We adapt such a scheme for the task of relighting and show our architecture in Fig. 2. Since we use a latent diffusion model (LDM), we first map all images and environment maps into the latent space of the original 2D diffusion model. In particular, the environment map is encoded in a similar manner to Neural Gaffer [6]: We use two separate latents corresponding to high dynamic range (HDR) and low dynamic range (LDR). This representation captures both bright details like direct light sources, as well as relatively dim sources like diffuse objects. Like Neural Gaffer, we also use standard tone mapping for the LDR environment map [38, 39], and for the HDR encoding we apply logarithmic tone mapping followed by normalization to $[0,1]$ by subtracting the minimal value and dividing by the maximal value. Additionally, for each image $I_{i}$ in our dataset, we apply a 3D rotation to the environment map to align it with the corresponding camera pose $\pi_{i}$. We combine the HDR and LDR encodings of the environment map with the encoded image and the ray map of the pose of each image by concatenating them, and then apply self-attention. Details of the network architecture for our multi-view relighting diffusion models are provided in Fig. 1 of the supplementary material.
@@ 3.2 Relightable Neural Radiance Field
We then use the multi-illumination dataset $\mathcal{D}_{\text{relit}}$ to train a NeRF. Since we wish to fit the NeRF model to our dataset with varying lighting so that it generalizes to novel lighting environments outside of our training set of illuminants $\mathcal{E}$, care must be taken when designing the model’s architecture. We use NeRF-Casting [40] as the base NeRF model since it efficiently captures view-dependent appearance, and we modify it to allow for conditioning on the illumination. Crucially, we encode the environment maps using two types of conditioning signals: general conditioning and specular conditioning. The general conditioning encodes the entire environment map into an embedding that is fed to the appearance MLP and is designed as a general-purpose, low-frequency signal for relighting, while the specular conditioning only encodes incident light coming from the specular direction. Its goal is to improve the model’s capacity to capture high-frequency reflections (similar to prior work on reconstructing specular reflections in NeRF [41]). The conditioning signals are illustrated in Fig. 3.
In order to encode the specular appearance of a camera ray, NeRF-Casting [40] casts a small set of reflected rays, traces them through the NeRF’s geometry, and volume renders a feature vector $\bar{\mathbf{f}}$ that encodes the scene content observed by these reflected rays. In our setting, we provide the two conditioning signals by concatenating them to the feature vector $\bar{\mathbf{f}}$ of each ray, which is mapped by NeRF-Casting’s MLP to the ray’s specular color component.
@@
Our general conditioning signal encodes the full environment map, and is therefore a complete description of the lighting of the object. To do this, we train an transformer-based encoder which maps the environment map into a vector. While the idea of using a per-lighting encoding is similar to the approach of NeRF-in-the-Wild [42], a crucial difference is that our embeddings are not optimized to fit to individual images, but they are instead parameterized as a learned mapping from the environment maps. This enables us to render the scene under novel illumination at inference time without requiring any additional training.
To encode the illumination features from the environment map, we utilize a transformer encoder with self-attention. Our transformer is trained from scratch, and its architecture is based on ViT-S8 of DINO [43], followed by a single matrix multiplication $W$ to produce a compact 128-dimensional embedding vector:
$\mathbf{f}^{\text{general}}=W\cdot\text{ViT}(E).$ (3)
@@
Although the general conditioning vector theoretically contains all information necessary for relighting, we found it to be insufficient for reconstructing and rendering high-frequency specular highlights. Our specular conditioning is designed to fix that by explicitly encoding incident light from the specular reflection direction, similar to the encoding used in prior work for reconstructing reflective objects and scenes [41, 40, 44]. Instead of only sampling the environment map value at the reflection direction, we use a series of blur kernels centered around the reflection direction to simulate the effects of materials with different roughnesses. The $i$-th component of this conditioning vector can be written as:
$\mathbf{f}_{i}^{\text{specular}}=\int_{\mathbb{S}^{2}}E(\boldsymbol{\omega}^{\prime})\text{G}(\boldsymbol{\omega}^{\prime};\boldsymbol{\omega}_{r},\sigma_{i})d\boldsymbol{\omega}^{\prime},$ (4)
@@
Network Predictions. Following NeRF-Casting [40], our relightable NeRF takes as input the 3D coordinates, ray direction, general conditioning, and specular conditioning. The geometry MLP predicts density, roughness, normals, and geometry features, while the color MLP outputs the RGB values.
@@
Figure 4: Qualitative comparisons on TensoIR [45]. All renderings are rescaled to the image resolution of the ground truth. Compared to previous works, our method recovers more plausible specular highlights and more accurate colors as indicated with the red box.
@@ 4.1 Datasets
To train multi-view relighting diffusion, we use a dataset of 400k synthetic 3D objects, including 100k from Objaverse [46]. Each object is rendered in 64 views $\times$ 16 HDR illuminations, with environments sampled from Polyhaven [47] (590 maps) and augmented via random up-axis rotations.
Evaluation datasets and metrics.
For relighting evaluations, we used two datasets: TensoIR [45] and Stanford-ORB [48]. TensoIR is a synthetic benchmark, which contains renderings of four objects under six lighting conditions. Stanford-ORB is real-world benchmarks by data capture in the wild. It has 14 objects composed of various materials. For quantitative comparisons, We evaluate the relighting rendering quality by PSNR, SSIM [49], and LPIPS-VGG [50].
@@ 4.2 Implementation details
We fine-tune our model starting from a pre-trained latent image generation model, as described in [51]. The multi-view denoiser is derived from the CAT3D network architecture [37], with modifications to the input channel dimensions to align with our relighting task. The inputs to our model are images with resolution $512\times 512$ which are encoded to $64\times 64\times 8$ latents. We chose the number of views to be 64. The model was trained on 128 TPU v5 chips using a learning rate of $10^{-4}$, with a total batch size of 128 for 360k iterations. After training, we generate the multi-illumination dataset by running our relighting diffusion inference on 111 environment maps.
@@ 4.3 Baselines
We compare our method against existing inverse rendering methods including NeRFactor [52], InvRender [53], PhySG [54], NeRD [19], NVDiffRecMC [14], NVDiffRec [15], Neural-PBIR [55], NeRO [44],TensoIR [45], and recent single-view relighting diffusion methods IllumiNeRF [7] and Neural Gaffer [6]. We also include the most recent Gaussian splatting-based inverse rendering method R3DG [56].
@@
NeRFactor [52] 23.38 0.908 0.131
InvRender [53] 23.97 0.901 0.101
TensoIR [45] 28.58 0.944 0.081
Neural-PBIR [55] 27.09 0.925 0.085
NeRO [44] 27.00 0.935 0.074
R3DG [56] 29.05 0.937 0.080
NeuralGaffer [6] 27.30 0.918 0.122
IllumiNeRF [7] 29.71 0.947 0.072
@@
Table 1: TensoIR benchmark [45]. We evaluate all four objects in the benchmark, each under five novel lightings. Each object is rendered from 200 views for novel view evaluation under each lighting. Thus, we have 4,000 renderings in total for quantitative evaluation. Best and 2nd-best are highlighted.

### v3

@@ 1 Introduction
Inserting real-world objects into new environments is a long-standing problem in computer graphics [debevec2008rendering, debevec2000acquiring], with numerous applications in the movie and gaming industries. While recent years have seen significant progress in 3D object reconstruction for view synthesis using radiance fields [Nerf, GaussianSplatting], these techniques represent an object illuminated by a single fixed environment and they do not enable changing the appearance of the object due to changes in the lighting. This work extends 3D object reconstruction to enable rendering the object under arbitrary illumination.
@@
At the same time, recent relighting diffusion models [zeng2024dilightnet, jin2024neural_gaffer, zhao2024illuminerf] have demonstrated impressive capabilities for generating realistic images of objects under arbitrary illumination. However, they only generate a single relit image at a time, which results in inconsistent relighting results when applied to a sequence of viewpoints. While these inconsistently-relit samples can be reconstructed into a single 3D representation [zhao2024illuminerf], optimizing a new 3D representation for each new target lighting is tedious and precludes interactive use cases.
@@ 2 Related Work
Recovering relightable 3D representations of objects from images is a longstanding goal in computer vision and graphics. The prevalent approach is inverse rendering: decomposing the object’s appearance into its underlying geometry, illumination, and material parameters, and relighting the object by simulating the physical interaction of a new target illumination with the recovered geometry and materials [ramamoorthi01, yu99, sato1997object, lensch2003image].
Modern methods for reconstructing relightable 3D representations with inverse rendering use differentiable rendering techniques [Li:2018:DMC, NimierDavidVicini2019Mitsuba2] to optimize mesh [hasselgren2022nvdiffrecmc, NVDiffRec], distance field [iron-2022], or volumetric [nrf, nerv2021, boss2021nerd, nerfactor, mai2023neural, tensoir, attal2024flash, feng2024arm] representations of object geometry and material parameters. While these inverse rendering approaches can be effective, they come with significant limitations: errors in estimating an object’s geometry and materials can produce unrealistic appearance when the object is relit under a new illumination and physically-accurate relighting involves computationally-expensive Monte Carlo simulation of global illumination, which can be too slow for interactive use cases.
@@
Early approaches for interactive relighting in the field of computer graphics proposed the idea of Precomputed Radiance Transfer (PRT) [sloan02]. As appearance behaves linearly with respect to lighting, an object or scene’s appearance can be precomputed under a set of basis lightings and these can be linearly combined to produce appearance under any desired target lighting condition. While PRT-based techniques enable interactive relighting, they struggle with the memory demands of storing the precomputed transfer matrices. To enable rendering under novel illuminations from arbitrary viewpoints, PRT-based methods need to store a full transfer matrix for each 3D point in the scene. These memory requirements are prohibitive even for modestly sized scenes and environment maps, so a large body of work focuses on compressing these PRT matrices [ng03, sloan03].
One-light-at-a-time (OLAT) captures in light stages [debevec00] can be thought of as directly capturing basis vectors of a real object’s radiance transfer matrices. Since each image captured by the light stage is of the object illuminated by a single element of a standard lighting basis, they can be linearly combined to reproduce the object’s appearance under any desired environment illumination.
@@
In the deep learning era, many methods have trained networks to directly output relit images. Early techniques utilized standard convolutional neural networks [xu2018deep, sun2019single] and more recent approaches are based on powerful diffusion models [zeng2024dilightnet, zhao2024illuminerf, jin2024neural_gaffer, poirier2024diffusion]. These generative relighting models have produced impressive results for image relighting, but they cannot be easily used for 3D relighting as the relit appearance is often not consistent across views. IllumiNeRF [zhao2024illuminerf], MIS [poirier2024diffusion], and Neural Gaffer [jin2024neural_gaffer] perform 3D relighting by reconciling samples from a 2D image relighting diffusion model into a single NeRF. However, the relit 3D representation needs to be optimized separately for each novel target illumination. Recent work on 3D reconstruction from images of an object with varying illumination [alzayer2024generative] uses a multi-view diffusion model to relight them to have consistent illumination. However, this approach does not allow relighting using a new unseen illumination condition, so it does not allow for generalization. And it does not utilize the high-frequency details in environment maps. Zeng et al. [zeng2023relighting] and RNG [fan2024rng] utilize shadow and highlight cues as conditioning signals for NeRF. In contrast, we explicitly use the encoded features of incident light from the specular reflection direction. Additionally, in contrast to all of these approaches, our proposed method trains a generalizable relightable NeRF that can be rendered under any target illumination without additional optimization. A concurrent work, RelitLRM [zhang2024relitlrm], directly recovers 3D geometry and appearance but does not guarantee consistent geometry across environment maps. In contrast, our method reconstructs constant geometry under varying illumination. DiffusionRenderer [liang2025diffusion] employs a video diffusion model for relighting but lacks 3D consistency, while our approach enforces it by explicitly modeling geometry and rendering. Moreover, our model enables fast inference, avoiding the long sampling times typical of diffusion-based generation.
@@ 3.1 Multi-View Diffusion Relighting
Our work crucially relies on a multi-view diffusion relighting model to provide consistent relit images, which can then be distilled into a relightable 3D model. This is in contrast with prior work on relighting using diffusion models [zhao2024illuminerf, jin2024neural_gaffer] which independently relights single images and results in ambiguities that must be resolved at the 3D reconstruction stage.
Our network architecture is inspired by multi-view diffusion architectures such as CAT3D [gao2024cat3d], which start with pretrained 2D image diffusion models and inflate them by adding cross-attention layers to process multiple views. We adapt such a scheme for the task of relighting and show our architecture in Fig. 2. Since we use a latent diffusion model (LDM), we first map all images and environment maps into the latent space of the original 2D diffusion model. In particular, the environment map is encoded in a similar manner to Neural Gaffer [jin2024neural_gaffer]: We use two separate latents corresponding to high dynamic range (HDR) and low dynamic range (LDR). This representation captures both bright details like direct light sources, as well as relatively dim sources like diffuse objects. Like Neural Gaffer, we also use standard tone mapping for the LDR environment map [hasinoff2016burst, debevec2004high], and for the HDR encoding we apply logarithmic tone mapping followed by normalization to $[0,1]$ by subtracting the minimal value and dividing by the maximal value. Additionally, for each image $I_{i}$ in our dataset, we apply a 3D rotation to the environment map to align it with the corresponding camera pose $\pi_{i}$. We combine the HDR and LDR encodings of the environment map with the encoded image and the ray map of the pose of each image by concatenating them, and then apply self-attention. Details of the network architecture for our multi-view relighting diffusion models are provided in Fig. 1 of the supplementary material.
@@ 3.2 Relightable Neural Radiance Field
We then use the multi-illumination dataset $\mathcal{D}_{\text{relit}}$ to train a NeRF. Since we wish to fit the NeRF model to our dataset with varying lighting so that it generalizes to novel lighting environments outside of our training set of illuminants $\mathcal{E}$, care must be taken when designing the model’s architecture. We use NeRF-Casting [verbin2024nerf] as the base NeRF model since it efficiently captures view-dependent appearance, and we modify it to allow for conditioning on the illumination. Crucially, we encode the environment maps using two types of conditioning signals: general conditioning and specular conditioning. The general conditioning encodes the entire environment map into an embedding that is fed to the appearance MLP and is designed as a general-purpose, low-frequency signal for relighting, while the specular conditioning only encodes incident light coming from the specular direction. Its goal is to improve the model’s capacity to capture high-frequency reflections (similar to prior work on reconstructing specular reflections in NeRF [verbin2022refnerf]). The conditioning signals are illustrated in Fig. 3.
In order to encode the specular appearance of a camera ray, NeRF-Casting [verbin2024nerf] casts a small set of reflected rays, traces them through the NeRF’s geometry, and volume renders a feature vector $\bar{\mathbf{f}}$ that encodes the scene content observed by these reflected rays. In our setting, we provide the two conditioning signals by concatenating them to the feature vector $\bar{\mathbf{f}}$ of each ray, which is mapped by NeRF-Casting’s MLP to the ray’s specular color component.
@@
Our general conditioning signal encodes the full environment map, and is therefore a complete description of the lighting of the object. To do this, we train an transformer-based encoder which maps the environment map into a vector. While the idea of using a per-lighting encoding is similar to the approach of NeRF-in-the-Wild [martinbrualla2020nerfw], a crucial difference is that our embeddings are not optimized to fit to individual images, but they are instead parameterized as a learned mapping from the environment maps. This enables us to render the scene under novel illumination at inference time without requiring any additional training.
To encode the illumination features from the environment map, we utilize a transformer encoder with self-attention. Our transformer is trained from scratch, and its architecture is based on ViT-S8 of DINO [DINO], followed by a single matrix multiplication $W$ to produce a compact 128-dimensional embedding vector:
$\mathbf{f}^{\text{general}}=W\cdot\text{ViT}(E)$ (3)
@@
Although the general conditioning vector theoretically contains all information necessary for relighting, we found it to be insufficient for reconstructing and rendering high-frequency specular highlights. Our specular conditioning is designed to fix that by explicitly encoding incident light from the specular reflection direction, similar to the encoding used in prior work for reconstructing reflective objects and scenes [verbin2022refnerf, verbin2024nerf, liu2023nero]. Instead of only sampling the environment map value at the reflection direction, we use a series of blur kernels centered around the reflection direction to simulate the effects of materials with different roughnesses. The $i$-th component of this conditioning vector can be written as:
$\mathbf{f}_{i}^{\text{specular}}=\int_{\mathbb{S}^{2}}E(\boldsymbol{\omega}^{\prime})\text{G}(\boldsymbol{\omega}^{\prime};\boldsymbol{\omega}_{r},\sigma_{i})d\boldsymbol{\omega}^{\prime}$ (4)
@@
Network Predictions. Following NeRF-Casting [verbin2024nerf], our relightable NeRF takes as input the 3D coordinates, ray direction, general conditioning, and specular conditioning. The geometry MLP predicts density, roughness, normals, and geometry features, while the color MLP outputs the RGB values.
@@
Figure 4: Qualitative comparisons on TensoIR [jin2023tensoir]. All renderings are rescaled to the image resolution of the ground truth. Compared to previous works, our method recovers more plausible specular highlights and more accurate colors as indicated with the red box.
@@ 4.1 Datasets
To train multi-view relighting diffusion, we use a dataset of 400k synthetic 3D objects, including 100k from Objaverse [deitke2023objaverse]. Each object is rendered in 64 views $\times$ 16 HDR illuminations, with environments sampled from Polyhaven [zaal2021polyhaven] (590 maps) and augmented via random up-axis rotations.
Evaluation datasets.
For relighting evaluations, we used two datasets: TensoIR [jin2023tensoir] and Stanford-ORB [kuang2024stanford]. TensoIR is a synthetic benchmark, which contains renderings of four objects under six lighting conditions. We use the train split of 100 views with “sunset” lighting condition as inputs for relightable NeRF. We then evaluate 200 novel views under other five environment maps, including “bridge”, “fireplace”, “forest”, “city”, and “night”. In total, we have 4,000 renderings for evaluation metric calculation. Stanford-ORB is real-world benchmarks by data capture in the wild. It has 14 objects composed of various materials. Each object is captured under three distinct lighting conditions, producing a total of 42 (object, lighting) combinations. Following its evaluation protocol, we use images of an object under a single lighting condition and evaluate novel views under the two target lighting settings.
@@ 4.2 Implementation details
We fine-tune our model starting from a pre-trained latent image generation model, as described in [rombach2022high]. The multi-view denoiser is derived from the CAT3D network architecture [gao2024cat3d], with modifications to the input channel dimensions to align with our relighting task. The inputs to our model are images with resolution $512\times 512$ which are encoded to $64\times 64\times 8$ latents. We chose the number of views to be 64. The model was trained on 128 TPU v5 chips using a learning rate of $10^{-4}$, with a total batch size of 128 for 360k iterations. After training, we generate the multi-illumination dataset by running our relighting diffusion inference on 111 environment maps.
@@ 4.3 Baselines
We compare our method against existing inverse rendering methods including NeRFactor [zhang2021nerfactor], InvRender [zhang2022modeling], PhySG [zhang2021physg], NeRD [boss2021nerd], NVDiffRecMC [hasselgren2022nvdiffrecmc], NVDiffRec [NVDiffRec], Neural-PBIR [Neural-PBIR], NeRO [liu2023nero],TensoIR [jin2023tensoir], and recent single-view relighting diffusion methods IllumiNeRF [zhao2024illuminerf] and Neural Gaffer [jin2024neural_gaffer]. We also include the most recent Gaussian splatting-based inverse rendering method R3DG [R3DG].
4.4 Evaluation Metrics
We evaluate the relighting rendering quality by PSNR, SSIM [SSIM], and LPIPS-VGG [LPIPS].
@@ 4.4 Evaluation Metrics
NeRFactor [zhang2021nerfactor] 23.38 0.908 0.131
InvRender [zhang2022modeling] 23.97 0.901 0.101
TensoIR [jin2023tensoir] 28.58 0.944 0.081
Neural-PBIR [Neural-PBIR] 27.09 0.925 0.085
NeRO [liu2023nero] 27.00 0.935 0.074
R3DG [R3DG] 29.05 0.937 0.080
NeuralGaffer [jin2024neural_gaffer] 27.30 0.918 0.122
IllumiNeRF [zhao2024illuminerf] 29.71 0.947 0.072
@@
Table 1: TensoIR benchmark [jin2023tensoir]. We evaluate all four objects in the benchmark, each under five novel lightings. Each object is rendered from 200 views for novel view evaluation under each lighting. Thus, we have 4,000 renderings in total for quantitative evaluation. Best and 2nd-best are highlighted.